table below shows the distribution of smokers and non – smokers by gender.

  Smokers Non – smokers Men 88 42 women 8 70

If one person is chosen at random find the probability that the person is:

A woman who smokes?

A non – smoker given that she is a woman

A man or a smoker.

(2 marks) 



1
Expert's answer
2021-11-26T05:21:14-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Smokers & Non – smokers & Total \\ \hline
 Men & 88 & 42 & 130 \\
 \hdashline
 Women & 8 & 70 & 78\\
\hdashline
 Total & 96 & 112 & 208\\
\end{array}"

1.

"P(W \& \ Smoker)=\dfrac{8}{208}=\dfrac{1}{26}"

2.


"P(Non-smoker|W)=\dfrac{70}{78}=\dfrac{35}{39}"

3.


"P(M\ or\ Smoker)=\dfrac{130+96-88}{208}=\dfrac{69}{104}"